# MalMixer: Few-Shot Malware Classification with Retrieval-Augmented Semi-Supervised Learning

### v2

@@ Abstract
Recent growth and proliferation of malware has tested practitioners’ ability to promptly classify new samples according to malware families. In contrast to labor-intensive reverse engineering efforts, machine learning approaches have demonstrated increased speed and accuracy. However, most existing deep-learning malware family classifiers must be calibrated using a large number of samples that are painstakingly manually analyzed before training. Furthermore, as novel malware samples arise that are beyond the scope of the training set, additional reverse engineering effort must be employed to update the training set. The sheer volume of new samples found in the wild creates substantial pressure on practitioners’ ability to reverse engineer enough malware to adequately train modern classifiers. In this paper, we present MalMixer, a malware family classifier using semi-supervised learning that achieves high accuracy with sparse training data. We present a novel domain-knowledge-aware technique for augmenting malware feature representations, enhancing few-shot performance of semi-supervised malware family classification. We show that MalMixer achieves state-of-the-art performance in few-shot malware family classification settings. Our research confirms the feasibility and effectiveness of lightweight, domain-knowledge-aware feature augmentation methods and highlights the capabilities of similar semi-supervised classifiers in addressing malware classification issues.
I Introduction
Malware has emerged as a paramount concern threatening software security, networks, and users. In 2021, approximately $15.45\%$ of computers worldwide faced at least one malware attack [1] and the number of new malware variants continue to grow at an alarming rate. For instance, compared to 2019, Windows business malware threats in 2021 increased by $85\%$ and consumer threats increased by $47\%$ [2]. One important approach to describe malware behavior and consequences is classification into distinct families (e.g., trojan, adware, etc.). Thus, prompt and accurate classification of malware is critical for effective defense against potential breaches.
@@ I Introduction
Thus, in the face of new, adaptive malware samples or zero-day threats, conventional malware family classifiers must be updated to handle new samples. Unfortunately, incorporating new malware samples requires substantial reverse engineering effort, which is both costly and time-consuming [11, 12, 13, 14], taking engineers with years of experiences hours to weeks to fully understand a single malware sample [15, 10]. Thus, malware analyzers and machine learning solutions degrade in performance over time as new malware samples appear in the wild, requiring incremental retraining [16, 17]. This again requires effort from sufficiently-trained experts to analyze new samples to provide a basis for such retraining.
@@
To effectively use a small, promptly-obtained labeled dataset, one key challenge is that a small dataset lacks the diversity seen in the wild and may not be representative of the entire cohort of related malware samples [20]. Thus, we desire techniques that enable robust and accurate classification of novel samples based on a relatively small corpus of manually-labeled malware samples — that is, a few shot approach to robust malware family classification.
@@
Previously, several papers have explored few-shot semi-supervised malware classifiers. For instance, CAD-FSL [21] presents a framework using code-aware GANs to synthesize variants of labeled malware binaries, enhancing a CNN classifier. Similarly, Catak et al. [22] employ image augmentation techniques like noise injection for generating new malware data. These works represent a popular semi-supervised framework that uses data augmentation to grow the size of input samples and enhance a downstream classifier. As data augmentation is well-established in computer vision [20], almost all existing work employing augmentation first converts malware binaries to grayscale images (byteplots) on which to employ image augmentation. That said, image-based techniques do not modify malware feature representations in a semantically-meaningful manner — informally, varying pixel data does not correspond to sensible malware properties. Other semi-supervsied classifiers [5, 23] often require a large amount of labeled data or employ simpler datasets with few malware families, limiting their real-world applicability. Thus, there is a need for augmentation techniques that more sensibly augment malware data representations by leveraging domain knowledge on the data’s semantics, and thereby demonstrating proficiency in practical scenarios.
@@
In this paper, we present MalMixer, a few-shot malware classifier unifying (1) a retrieval-based, domain-knowledge-aware malware feature augmentation method that is aware of the underlying semantics of each malware feature, and (2) a semi-supervised learning framework that accentuates the effect of our augmentation approach. This combination provides robust and accurate family classification with few labeled malware samples. At a high level, we grow a malware dataset through domain-knowledge-aware data augmentation on malware features. As shown in Figure 1, we divide malware features into interpolatable and non-interpolatable sets (see our proposed definitions in Section III-B), treating them differently during augmentation. Then, we use a similarity-based retrieval technique to mutate non-interpolatable malware features and use manifold alignment to maximize the compatibility between the augmented interpolatable and non-interpolatable features. Finally, the augmented data is used to enhance a semi-supervised framework that combines consistency regularization, entropy minimization, and traditional regularization, which follows best practices from other domains [24]. With these components, our classifier enables practitioners to efficiently and accurately classify large malware datasets by requiring only a few labeled samples of families of interest, dramatically improving the prediction process.
@@
Figure 1: Overview of our malware augmentation approach. We divide static malware features into interpolatable and non-interpretable sets, and use retrieval and alignment techniques to generate synthetic samples.
@@
    We design a novel similarity-and-retrieval-based augmentation technique on malware feature representations to generate synthetic data.
@@
    We derive a manifold alignment technique to encourage consistency between all synthesized feature values. As such, our augmentation generates synthetic samples that more closely mimic ground-truth family distributions.
@@
    We conduct extensive experiments on benchmark datasets in evaluation, demonstrating that MalMixer outperforms other state-of-the-art static-feature-based malware classifiers in multiple few-shot settings.
II Background and Related Work
In this section, we highlight the areas of (1) malware analysis methods to generate feature representations of malware, (2) augmentation methods that grow training sets, and the landscape of (3) semi-supervised learning and (4) domain-invariant learning, from which we derive our classifier model.
@@ II Background and Related Work
Malware Data Augmentation Data augmentation is a technique used to expand a dataset by applying input transformations that do not alter the underlying class semantics [24]. Well-established within computer vision communities, conventional augmentation uses mutations such as image zooming, cropping, and noise injection [27] to create novel samples that can be used to further train an underlying model. Many security researchers have explored converting malware binaries into grayscale images, and subsequently applying image-based augmentation techniques [28, 29, 30, 21, 22]. Nevertheless, these image-based techniques are rather noise-inducing rather than semantically meaningful to malware features. Recently, Wong et al. [20] demonstrated the feasibility of semantically-meaningful augmentation of malware binaries. The researchers performed semantics-preserving code transformations on labeled binaries to generate new mutated malware samples, therefore boosting the performance of ML-driven classifiers. However, this is still time-consuming and resource-intensive and can only be practically conducted on a small number of samples [20]. In this paper, we present a lightweight feature-based, domain-knowledge-aware augmentation technique, guided by the semantics of the static malware features.
@@
Semi-Supervised Learning Semi-supervised learning is widely used to tackle data scarcity issues by combining limited labeled data with a volume of unlabeled data during the training process. Popular strategies for semi-supervised learning progressively update the unlabeled dataset with high-confidence labeled data using techniques like label mixing and bootstrapping [31, 32], curriculum learning [33, 34], and meta-learning [35, 36]. Concurrently, they also aim to introduce regularization terms to ensure model consistency  [37, 24, 23]. In malware classification, several studies have employed semi-supervised learning techniques, such as malware embedding [23, 11], transfer learning [38], and soft labeling [39]. MalMixer elaborates upon former works by proposing a novel data-centric augmentation method to enhance semi-supervised learning.
@@
Domain Invariant Learning Domain Invariant Learning is a subfield of machine learning focused on generalizing models across different but related domains. Researchers aim to identify invariant features in data that characterize common trends amongst different domains that assist downstream model performance. One significant category of Domain Invariant Learning is feature invariance learning, in which domain shift is reduced using unsupervised techniques to extract domain-invariant features [40, 41]. In this paper, we use feature invariance learning to identify common malware features across malware families to better generate synthetic samples as part of a malware augmentation pipeline.
III Approach
@@
In this section, we present MalMixer, a few-shot malware family classifier unifying retrieval-based augmentation and semi-supervised learning. We also introduce the underlying assumptions we make in constructing this approach.
@@
Recall the goal of our approach is to identify the malware family of a given input malware sample. To achieve this, we train a classifier on a large dataset, with only a small subset of malware samples well-labeled. For each sample in the dataset, we apply static analyses to extract features. Then, within the feature space, we apply our augmentation technique to each original sample in the dataset to generate more novel samples that closely reflect the distributions of the original samples. Finally, we feed all augmented malware samples into a semi-supervised classification framework that accentuates the effect of our proposed data augmentation.
@@
Before proceeding to model training, each malware executable is statically analyzed (Section V-A) to extract features. As such, MalMixer is primarily geared towards enhancing static-feature-based malware classification. This domain is important for several reasons: (1) extracted features typically yield superior classification results compared to end-to-end classifiers that process entire malware binaries [42], (2) static features can be promptly derived from new datasets, and (3) due to the rise of sophisticated sandbox-evasion techniques, there is a consistent need to shift back to static analysis in industrial settings [43]. Typical static analysis features include file size, hash of imported libraries, target OS, section information, byte entropy, and others [25]. In general, static features are not engineered for detecting certain families but rather provide a comprehensive summary of a PE file. We evaluate MalMixer on a set of state-of-the-art, statically extracted PE feature representations in the EMBER format [42].
@@
Figure 2 illustrates MalMixer, which consists of (1) a retrieval-based augmentation pipeline, and (2) an overarching semi-supervised classification framework that accentuates the effect of the augmentation pipeline.
III-A Semi-Supervised Framework
@@
First, we present MalMixer’s backbone: an overarching, augmentation-based semi-supervised classification framework. By choosing semi-supervised learning, we seek to build a malware classifier that requires only a small portion of the dataset to be labeled. In practice, our approach has the potential to reduce the expensive manual effort in reverse engineering and labeling malware samples.
@@
Our framework is developed based on MixMatch [24], a semi-supervised learning framework originally designed for image classification. MixMatch is designed to transfer knowledge preserved in minimal labeling to largely unexplored domains. Thus, we adapt the MixMatch architecture for malware family classification, allowing us to expand a small corpus of labeled malware to a much larger, unlabeled dataset. MixMatch excels with semantics-preserving augmentation, but traditional image-based augmentation techniques (e.g., adding noise to byteplots [44]) do not respect malware samples’ semantics (Section II). Thus, MalMixer innovates upon prior art via domain-knowledge-aware augmentation tailored for malware data augmentation by considering malware features’ underlying semantics. The semi-supervised framework and our proposed augmentation approach together can achieve outstanding classification accuracy and resiliency against unseen malware samples, as showcased in Section VII.
@@
We now introduce the overarching semi-supervised framework adopted from MixMatch. At a high level, unlabeled malware is combined with ground truth labeled malware through data augmentation, label guessing, averaging, sharpening, and MixUp [37], before training a base classifier.
@@
Data Augmentation We denote the set of labeled data as $\mathcal{X}$, their labels as $\mathcal{Y}$, and unlabeled data as $\mathcal{U}$. Note that $\mathcal{X}$ and $\mathcal{U}$ are feature representations of malware executables. First, we apply data augmentation to $\mathcal{X}$ and $\mathcal{U}$ and transform them into $\mathcal{X^{\prime}}$ and $\mathcal{U^{\prime}}$. Specifically, with $\cup$ denoting multiset union,
$\mathcal{X^{\prime}}=\mathcal{X}\cup\mathcal{A}(\mathcal{X}),~{}\mathcal{U}^{\prime}=\mathcal{U}\cup\mathcal{A}(\mathcal{U}),$ (1)
where $\mathcal{A}$ is any stochastic data augmentation process s.t.
@@
with $label$ being a function that provides the ground truth label for a malware sample. Our proposed augmentation pipeline will be introduced in subsection (B).
Label Guessing, Average, & Sharpening, MixUp Next, we predict the pseudo labels for the unlabeled dataset (both original and augmented). We average and sharpen [24] the predicted labels for each original unlabeled sample and its corresponding augmented synthetic sample. The steps above provide us pseudo labels for the unlabeled data. Combining the original labeled dataset and pseudo labels on the unlabeled dataset, we obtain a fully labeled dataset. Next, we apply MixUp [37] to this dataset, and the output of MixUp is used to train a base classifier (e.g. ResNet). The base classifier is then used to classify test sets. A detailed formulation of all steps above can be found in Section A in the Appendix.
Figure 2: Illustration of our model framework, which consists of (1) a retrieval-based augmentation pipeline that grows the number of training samples and (2) an overarching augmentation-based semi-supervised classification framework that bolsters the effect of the augmentation pipeline.
@@
III-B Retrieval-Based Augmentation Pipeline
Within the aforementioned semi-supervised framework, a key component is data augmentation, $\mathcal{A}$, used in Equation (1). In malware family classification, it is challenging to completely satisfy Equation (2) due to difficulties with analyzing and reasoning about raw binaries [20]. In this subsection, we introduce an augmentation pipeline that approximates the equality in Equation (2) using high-confidence, domain-knowledge-aware mutations of each malware instance, denoted $s$. The outcome is a set of synthetic malware samples that are well-grounded and reflect more realistic, plausible samples. The augmentation process is decoupled from our semi-supervised framework and can be performed prior to training.
Our proposed augmentation pipeline mutates a malware sample by interpolating its features with a similar sample likely to be of the same family. However, the key challenge is that straightforward interpolations are not necessarily semantically meaningful — for example, taking the average of two MD5 hash values of strings in malware samples will not provide a meaningful, novel value.
To address the challenge, we follow four steps: (i) separating malware features into interpolatable and non-interpolatable sets, (ii) employing domain-invariant learning to project static features of malware into embedding spaces, (iii) applying a retrieval-based method to identify candidate features representative of a mixture of two malware instances, (iv) using an alignment-based approach to generate the best synthetic malware sample whose interdependent features are most compatible. Steps (i) and (ii) generate the oracles necessary for steps (iii) and (iv) that carry out the actual augmentation.

### v4

@@ Abstract
Recent growth and proliferation of malware have tested practitioners’ ability to promptly classify new samples according to malware families. In contrast to labor-intensive reverse engineering efforts, machine learning approaches have demonstrated increased speed and accuracy. However, most existing deep-learning malware family classifiers must be calibrated using a large number of samples that are painstakingly manually analyzed before training. Furthermore, as novel malware samples arise that are beyond the scope of the training set, additional reverse engineering effort must be employed to update the training set. The sheer volume of new samples found in the wild creates substantial pressure on practitioners’ ability to reverse engineer enough malware to adequately train modern classifiers.
In this paper, we present MalMixer, a malware family classifier using semi-supervised learning that achieves high accuracy with sparse training data. We present a domain-knowledge-aware data augmentation technique for malware feature representations, enhancing few-shot performance of semi-supervised malware family classification. We show that MalMixer achieves state-of-the-art performance in few-shot malware family classification settings. Our research confirms the feasibility and effectiveness of lightweight, domain-knowledge-aware data augmentation methods for malware features and shows the capabilities of similar semi-supervised classifiers in addressing malware classification issues.
1 Introduction
Malware has emerged as a paramount concern threatening software security, networks, and users. In 2021, approximately $15.45\%$ of computers worldwide faced at least one malware attack [1] and the number of new malware variants continue to grow at an alarming rate. For instance, compared to 2019, Windows business malware threats in 2021 increased by $85\%$ and consumer threats increased by $47\%$ [2]. One important approach to describe malware behavior and consequences is classification into distinct families (e.g., sfone, wacatac, etc.). Thus, prompt and accurate classification of malware is critical for effective defense against potential breaches.
@@ 1 Introduction
In the face of new, adaptive malware samples or zero-day threats, conventional malware family classifiers must be updated to handle new samples. Unfortunately, incorporating new malware samples requires substantial reverse engineering effort, which is both costly and time-consuming [11, 12, 13, 14], taking engineers with years of experiences hours to weeks to fully understand a single malware sample [15, 10]. Thus, malware analyzers and machine learning solutions degrade in performance over time as new malware samples appear in the wild, requiring incremental retraining [16, 17]. This again requires effort from sufficiently-trained experts to analyze new samples to provide a basis for such retraining.
@@
To effectively use a small, promptly-obtained labeled dataset, one key challenge is that a small dataset may not be sufficiently diverse or representative of the entire cohort of related malware samples [20]. Thus, we desire techniques that enable robust and accurate classification of novel samples based on a relatively small corpus of manually-labeled malware samples — that is, a few-shot approach to robust malware family classification.
@@
Previously, several papers have explored few-shot semi-supervised malware classifiers. For instance, CAD-FSL [21] presents a framework using code-aware GANs to synthesize variants of labeled malware binaries, enhancing a CNN classifier. Similarly, Catak et al. [22] employ image augmentation techniques like noise injection for generating new malware data. These works use data augmentation to grow the size of input samples and enhance a downstream classifier. As data augmentation is well-established in computer vision [20], almost all existing work employing augmentation first converts malware binaries to grayscale images (byteplots) on which to employ image augmentation. That said, image-based techniques do not modify malware feature representations in a semantically-meaningful manner — informally, varying pixel data does not correspond to sensible malware properties. Compared to image-based techniques, semantics-preserving data augmentation methods exist, but typically require semantics-aware transformations on malware binaries [20], which can be very time and resource-intensive. Thus, there is a need for data augmentation techniques that more sensibly augment malware feature representations by leveraging domain knowledge on the data’s semantics, while also remaining computationally efficient for real-world applicability. In designing our approach, we perform semantics-aware data augmentation on malware’s feature representations, making a deliberate trade-off between some semantic validity and operational efficiency to achieve a practical solution.
@@
In this paper, we present MalMixer, a few-shot malware classifier unifying (1) a retrieval-based, domain-knowledge-aware malware feature augmentation method that is aware of the underlying semantics of each malware feature, and (2) a semi-supervised learning framework that accentuates the effect of our augmentation approach. This combination provides robust and accurate family classification with few labeled malware samples. At a high level, we grow a malware dataset through domain-knowledge-aware data augmentation on malware features. As shown in Figure 1, we divide malware features into interpolatable and non-interpolatable sets (defined in Section 3.2) that are treated differently during training. Then, we use a similarity-based retrieval technique to mutate non-interpolatable malware features and use manifold alignment to maximize the compatibility between the augmented interpolatable and non-interpolatable features. Finally, the augmented pseudo-malware samples are used to enhance a semi-supervised framework that combines consistency regularization, entropy minimization, and traditional regularization, following best practices in other domains [23]. The resulting approach enables efficiently and accurately classifying large malware datasets by requiring only a few labeled samples of families of interest.
@@
Figure 1: Overview of our proposed data augmentation approach for malware feature representations. We divide static malware features into interpolatable and non-interpolatable sets, and use retrieval and alignment techniques to generate synthetic samples.
@@
    We design a similarity-and-retrieval-based data augmentation technique on malware feature representations to generate synthetic data.
@@
    We derive a manifold alignment technique to encourage consistency between all synthesized feature values to synthesize samples that mimic ground-truth family distributions.
@@
    We thoroughly evaluate MalMixer, showing we outperform other state-of-the-art static-feature-based malware classifiers in few-shot settings.
2 Background and Related Work
In this section, we discuss (1) malware analysis methods to generate feature representations of malware, (2) data augmentation methods that grow training sets, (3) semi-supervised learning and (4) domain-invariant learning, from which we derive our classifier model.
@@ 2 Background and Related Work
Malware Data Augmentation Data augmentation is a technique used to expand a dataset by applying input transformations that do not alter the underlying class semantics [23]. Well-established within computer vision communities, conventional augmentation uses mutations such as image zooming, cropping, and noise injection [26] to create novel samples that can be used to further train an underlying model. Many security researchers have explored converting malware binaries into grayscale images, and subsequently applying image-based augmentation techniques [27, 28, 29, 21, 22]. Nevertheless, these image-based techniques are rather noise-inducing rather than semantically meaningful to malware features. Recently, Wong et al. [20] demonstrated the feasibility of semantically-meaningful data augmentation of malware binaries. The researchers performed semantics-preserving code transformations on labeled binaries to generate new mutated malware samples, therefore boosting the performance of ML-driven classifiers. However, this is still time-consuming and resource-intensive and can only be practically conducted on a small number of samples [20]. In this paper, we present a lightweight feature-based, domain-knowledge-aware data augmentation technique, guided by the semantics of the static malware features.
@@
Semi-Supervised Learning Semi-supervised learning is widely used to address data scarcity issues by combining limited labeled data with a volume of unlabeled data during the training process. Popular strategies for semi-supervised learning progressively update the unlabeled dataset with high-confidence labeled data using techniques like label mixing and bootstrapping [30, 31], curriculum learning [32, 33], and meta-learning [34, 35]. Concurrently, they also aim to introduce regularization terms to ensure model consistency [36, 23, 37]. In malware classification, several studies have employed semi-supervised learning techniques, such as malware embedding [37, 11], transfer learning [38], and soft labeling [39]. MalMixer builds upon former works by introducing a novel data-centric augmentation method to enhance semi-supervised learning.
@@
Domain Invariant Learning Domain Invariant Learning is a subfield of machine learning focused on generalizing models across different but related domains. Researchers aim to identify invariant features in data that characterize common trends amongst different domains that assist downstream model performance. One important category of Domain Invariant Learning is feature invariance learning, in which domain shift is reduced using unsupervised techniques to extract domain-invariant features [40, 41]. In this paper, we use feature invariance learning to identify common malware features across malware families to better generate synthetic samples as part of a malware data augmentation pipeline.
3 Approach
@@
We present MalMixer, a few-shot malware family classifier unifying retrieval-based data augmentation and semi-supervised learning. We also introduce the underlying assumptions we make in constructing this approach.
@@
Recall the goal of our approach is to identify the malware family of a given input malware sample. To achieve this, we train a classifier on a large dataset with only a small subset of expert-labeled malware samples. For each sample in the dataset, we apply static analyses to extract features. Then, within the feature space, we apply our data augmentation technique to each original sample in the dataset to generate more novel samples that closely reflect the distributions of the original samples. Finally, we feed all augmented pseudo-malware samples into a semi-supervised classification framework that accentuates the effect of our proposed data augmentation.
@@
Before proceeding to model training, each malware executable is statically analyzed (Section 5.1) to extract features. As such, MalMixer is primarily geared towards enhancing static-feature-based malware classification. This domain is important for several reasons: (1) extracted features typically yield superior classification results compared to end-to-end classifiers that process entire malware binaries [42], (2) static features can be promptly derived from new datasets, and (3) the rise of sophisticated sandbox-evasion techniques has led to a shift to static analysis in industrial settings [43]. Typical static analysis features include file size, hash of imported libraries, target OS, section information, byte entropy, among others [24]. In general, static features are not engineered for detecting certain families but rather provide a comprehensive summary of a PE file. We evaluate MalMixer on a set of state-of-the-art, statically extracted PE feature representations in the EMBER format [42].
@@
Figure 2 illustrates MalMixer, which consists of (1) a retrieval-based data augmentation pipeline, and (2) an overarching semi-supervised classification framework that accentuates the effect of the data augmentation pipeline.
Figure 2: Illustration of our model framework, which consists of (1) a retrieval-based augmentation pipeline that grows the number of training samples and (2) an overarching augmentation-based semi-supervised classification framework that bolsters the augmentation pipeline.
3.1 Semi-Supervised Framework
@@
First, we present MalMixer’s backbone: an data-augmentation-based semi-supervised classification framework. We use semi-supervised learning to build a malware classifier that requires only a small portion of the dataset to be labeled. In practice, our approach has the potential to reduce the expensive manual effort in reverse engineering and labeling malware samples.
@@
Our framework is based on MixMatch [23], a semi-supervised learning framework originally designed for image classification. MixMatch transfers knowledge preserved in minimal labeling to largely unexplored domains. Thus, we adapt MixMatch for malware family classification, allowing us to expand a small corpus of labeled malware to a much larger, unlabeled dataset. MixMatch excels with semantics-preserving data augmentation, but traditional image-based augmentation techniques (e.g., adding noise to byteplots [44]) do not respect malware samples’ semantics (Section 2). Thus, MalMixer innovates upon prior art via domain-knowledge-aware augmentation tailored for malware data augmentation by considering malware features’ underlying semantics. The semi-supervised framework and our augmentation approach can achieve outstanding classification accuracy and resiliency against unseen malware samples, as shown in Section 7.
@@
We now introduce the overarching semi-supervised framework adopted from MixMatch. At a high level, unlabeled malware is combined with ground truth labeled malware through data augmentation, label guessing, averaging, sharpening, and MixUp [36], before training a base classifier. In this paper, we use a Fully Connected ResNet (FC-ResNet) as our base classifier (Appendix A).
@@
Data Augmentation We denote the set of labeled data as $\mathcal{X}$, their labels as $\mathcal{Y}$, and unlabeled data as $\mathcal{U}$. Note that $\mathcal{X}$ and $\mathcal{U}$ are feature representations of malware executables. First, we apply data augmentation to $\mathcal{X}$ and $\mathcal{U}$ and transform them into $\mathcal{A}(\mathcal{X})$ and $\mathcal{A}(\mathcal{U})$, where $\mathcal{A}$ is any stochastic data augmentation process s.t.
@@
with $label$ being a function that provides the ground truth label for a malware sample. Our augmentation pipeline will be introduced in Subsection 3.2.
Label Guessing, Average, Sharpening, and MixUp Next, we generate pseudo labels for the unlabeled dataset. The base classifier $P$ predicts label probability distributions over all labels for both $\mathcal{U}$ and $\mathcal{A}(\mathcal{U})$. We take the average of these predictions to obtain more stable pseudo-label probabilities $\mathcal{Y_{\mathcal{U}}}=\frac{P(Y|\mathcal{U})+P(Y|\mathcal{A}(\mathcal{U}))}{2}$. Intuitively, this step encourages the classifier to assign consistent, similar labels to the original and augmented versions of the same malware.
We then sharpen pseudo-label probabilities $\mathcal{Y_{\mathcal{U}}}$ with $\text{Sharpen}(y)_{i}=\frac{y_{i}^{2}}{\sum_{j}y_{j}^{2}}$. This step is akin to applying softmax on pseudo-label distributions to reduce uncertainty by enhancing confident label predictions.
These steps provide us pseudo-label probabilities for each original unlabeled sample and its corresponding augmented synthetic sample. Combining the labeled dataset and the pseudo-labeled unlabeled dataset, we construct a fully labeled dataset $\mathcal{X^{\prime}}=(\mathcal{X}\cup\mathcal{A}(\mathcal{X}),\mathcal{Y}\cup\mathcal{Y})$ and $\mathcal{U^{\prime}}=(\mathcal{U}\cup\mathcal{A}(\mathcal{U}),\mathcal{Y_{\mathcal{U}}}\cup\mathcal{Y_{\mathcal{U}}})$.
Next, we apply MixUp [36] to labeled and pseudo-labeled samples. MixUp performs linear interpolation of features and label probabilities, giving us $\mathcal{X^{\prime}}=M(\mathcal{X^{\prime}},\mathcal{W_{X^{\prime}}}),~{}\mathcal{U^{\prime}}=M(\mathcal{U^{\prime}},\mathcal{W_{U^{\prime}}})$. Here, $M$ is linear interpolation and $\mathcal{W_{X^{\prime}}},\mathcal{W_{U^{\prime}}}$ are randomly selected subsets of the dataset, so that every sample is mixed up with some other sample. Intuitively, if feature representation $[0,0]$ has label probabilities $[1,0,0]$ and and feature $[2,2]$ has label $[0,1,0]$, MixUp encourages the classifier to assign label probabilities $[\frac{1}{2},\frac{1}{2},0]=M([1,0,0],[0,1,0])=\frac{1}{2}[1,0,0]+\frac{1}{2}[0,1,0]$ to feature $[1,1]=M([0,0],[2,2])=\frac{1}{2}[0,0]+\frac{1}{2}[2,2])$. This encourages the classifier to behave linearly between samples, smoothing decision boundaries and improving robustness.
To train a base classifier, the loss is
$\mathcal{L_{X^{\prime}}}$ $=\frac{1}{|\mathcal{X^{\prime}}|}\sum_{x,y\in\mathcal{X^{\prime}}}{\text{CrossEntropy}(y,P(y|x))},$
The final loss is $\mathcal{L}=\mathcal{L_{X^{\prime}}}+\mathcal{L_{U^{\prime}}}$. The trained base classifier $P$ is then used to classify test sets.
@@
3.2 Retrieval-Based Data Augmentation Pipeline
Within the aforementioned semi-supervised framework, a key component is data augmentation, $\mathcal{A}$, used in Equation (1). In malware family classification, it is challenging to completely satisfy Equation (1) due to difficulties with analyzing and reasoning about raw binaries [20]. In this subsection, we introduce a data augmentation pipeline that approximates the equality in Equation (1) using high-confidence, domain-knowledge-aware mutations of each malware instance, denoted $s$. The outcome is a set of synthetic malware samples that are well-grounded and reflect more realistic, plausible samples. The augmentation process is decoupled from our semi-supervised framework and can be performed prior to training.
Our proposed data augmentation pipeline mutates a malware sample by interpolating its features with a similar sample likely to be of the same family. However, the key challenge is straightforward interpolations are not necessarily semantically meaningful — for example, taking the average of two MD5 hash values of strings in malware samples will not provide a meaningful, novel value.
To address the challenge, we follow four steps: (i) separating malware features into interpolatable and non-interpolatable sets, (ii) employing domain-invariant learning to project static features of malware into embedding spaces, (iii) applying a retrieval-based method to identify candidate features representative of a mixture of two malware instances, (iv) using an alignment-based approach to generate the best synthetic malware sample whose interdependent features are most compatible. Steps (i) and (ii) generate the oracles necessary for steps (iii) and (iv) that carry out the actual data augmentation.